OpenStudy (anonymous):

simplify 3 square root of 125 75 square root of 5 15 square root of 5 8 square root of 5 5 square root of 5

OpenStudy (anonymous):

i need help like step by step and to be able to do a problem like this Simplify 3 fourth root of 112 48 fourth root of 7 2 fourth root of 7 5 fourth root of 7 6 fourth root of 7

OpenStudy (anonymous):

\[3 \sqrt{125} = 3 \sqrt{ 5 \times 5 \times 5} = 3 \times 5 \sqrt{5}\] How much??

OpenStudy (anonymous):

b